# PromptTSS: A Prompting-Based Approach for Interactive Multi-Granularity Time Series Segmentation

### v1

@@ Abstract.
Multivariate time series data, collected across various fields such as manufacturing and wearable technology, exhibit states at multiple levels of granularity, from coarse-grained system behaviors to fine-grained, detailed events. Effectively segmenting and integrating states across these different granularities is crucial for tasks like predictive maintenance and performance optimization. However, existing time series segmentation methods face two key challenges: (1) the inability to handle multiple levels of granularity within a unified model, and (2) limited adaptability to new, evolving patterns in dynamic environments. To address these challenges, we propose PromptTSS, a novel framework for time series segmentation with multi-granularity states. PromptTSS uses a unified model with a prompting mechanism that leverages label and boundary information to guide segmentation, capturing both coarse- and fine-grained patterns while adapting dynamically to unseen patterns. Experiments show PromptTSS improves accuracy by 24.49% in multi-granularity segmentation, 17.88% in single-granularity segmentation, and up to 599.24% in transfer learning, demonstrating its adaptability to hierarchical states and evolving time series dynamics.
@@
ccs: Mathematics of computing Time series analysisccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Supervised learning
@@ 5. Experiments
To create datasets with multiple levels of state granularity, we systematically merge adjacent states in the original fine-grained states. For Pump V35, Pump V36, and Pump V38, we generate both 2$\times$ and 4$\times$ coarser versions by progressively grouping consecutive states. For USC-HAD, we only create a 2$\times$ coarser version because the original dataset contains only 12 states, and further merging would result in an excessively small number of states, limiting meaningful segmentation analysis. We do not apply this state merging process to MoCap and ActRecTut, as their original number of states is already limited, and further merging would result in an insufficient number of states for robust analysis.
@@ 5.4.3. Sliding Window Length
The effect of sliding window length $T$ is evaluated in terms of both segmentation accuracy and training efficiency. As shown in Figure 6, extremely short windows (e.g., $T=128$) limit temporal context, leading to lower segmentation accuracy. Accuracy improves with larger window sizes as the model captures richer temporal information. PromptTSS maintains high accuracy even with large $T$, and achieves strong performance at $T=2048$ while incurring minimal additional training time. Thanks to its patching mechanism, training remains efficient across all window lengths, demonstrating that PromptTSS scales well to long input sequences.
@@ 5.4.4. Encoder Architecture
PromptTSS is evaluated with different backbone encoders: Transformer, ResNet, TCN, and LSTM. As shown in Figure 7, the Transformer consistently outperforms the other architectures on both USC-HAD and PumpV35. This result highlights its superior ability to model long-range temporal dependencies, justifying its selection as the default encoder for PromptTSS.

### v2

@@ Abstract.
Multivariate time series data, collected across various fields such as manufacturing and wearable technology, exhibit states at multiple levels of granularity, from coarse-grained system behaviors to fine-grained, detailed events. Effectively segmenting and integrating states across these different granularities is crucial for tasks like predictive maintenance and performance optimization. However, existing time series segmentation methods face two key challenges: (1) the inability to handle multiple levels of granularity within a unified model, and (2) limited adaptability to new, evolving patterns in dynamic environments. To address these challenges, we propose PromptTSS, a novel framework for time series segmentation with multi-granularity states. PromptTSS uses a unified model with a prompting mechanism that leverages label and boundary information to guide segmentation, capturing both coarse- and fine-grained patterns while adapting dynamically to unseen patterns. Experiments show PromptTSS improves accuracy by 24.49% in multi-granularity segmentation, 17.88% in single-granularity segmentation, and up to 599.24% in transfer learning, demonstrating its adaptability to hierarchical states and evolving time series dynamics. Our code is available at https://github.com/blacksnail789521/PromptTSS.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761142isbn: 979-8-4007-2040-6/2025/11ccs: Mathematics of computing Time series analysisccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Supervised learning
@@ 5. Experiments
[…]
@@ 5.4.3. Sliding Window Length
The effect of sliding window length $T$ is evaluated in terms of segmentation accuracy and training efficiency. As shown in Figure 6, extremely short windows (e.g., $T=128$) limit temporal context, leading to lower segmentation accuracy. Accuracy improves with larger window sizes as the model captures richer temporal information. PromptTSS maintains high accuracy even with large $T$, achieving strong performance at $T=2048$ while incurring minimal additional training time. Thanks to its patching mechanism, training remains efficient across all window lengths, demonstrating PromptTSS scales well to long input sequences.
@@ 5.4.4. Encoder Architecture
PromptTSS is evaluated with different backbone encoders: Transformer, ResNet, TCN, and LSTM. As shown in Figure 7, the Transformer consistently outperforms others on USC-HAD and PumpV35, demonstrating superior long-range temporal modeling and justifying its use as the default encoder.
@@ 6. Conclusion
Acknowledgements.
The authors would like to thank GoEdge.ai for providing internship support, computing resources, and valuable domain expertise that contributed to this research.
GenAI Usage Disclosure
In accordance with the ACM policy and CIKM 2025 submission guidelines, we disclose the following use of Generative AI in the preparation of this work: Generative AI tools (e.g., ChatGPT) were used solely for language polishing purposes, including minor grammar correction, sentence restructuring, and clarity improvements. No content, code, or experimental results were generated or influenced by these tools.